OpenStudy (anonymous):

Solve the system of equations: 1/x + 1/y = 8 3/x - 5/y = 0 The textbook gave me a hint that I'm supposed to use substitution. I tried that, but I couldn't get anywhere.

OpenStudy (ranga):

Let u = 1/x and v = 1/y Make the substitutions. Solve for u and v then put back u = 1/x; v = 1/y. Take the reciprocal to get x and y.

OpenStudy (anonymous):

Yeah, that's what the books says, but I don't really know how that works. Should I be getting u = 8 - v and v = 8-u?

OpenStudy (ranga):

After the substitution, you get 2 equations: u + v = 8 3u - 5v = 0 Solve for u and v. Use elimination method or substitution method -- whichever is easy for you.

OpenStudy (anonymous):

Ok! When I solved for u and v I got u=20 and v=-12. So from there I plug it back into u=1/x and v=1/y? 20=1/x => x=1/20 12=1/y => y=-1/12 I know I'm doing something wrong because the back says it's (1/5, 1/3)

OpenStudy (ranga):

You got the wrong values for u and v. u + v = 8 3u - 5v = 0 To eliminate v, multiply first equation by 5 and add it to the second: 5u + 5v = 40 3u - 5v = 0 add 8u = 40 divide both sides by 8 u = 5 put this back in first equation: 5 + v = 8 subtract 5 from both sides v = 8 - 5 = 3 u = 5; v = 3 put back u = 1/x and v = 1/y 1/x = 5 take reciprocal on both sides: x = 1/5 1/y = 3 y = 1/3

OpenStudy (ranga):

write it as (x,y): (1/5, 1/3) You may want to get some practice solving 2 equations and 2 unknowns because you will come across it many times in algebra.

OpenStudy (anonymous):

Oh! Thank you! I didn't notice that I wrote down 3u + 5v instead of 3u - 5v. This was really the only problem that confused me; I just need to keep watching out for careless mistakes. Thanks again!

OpenStudy (ranga):

you are welcome.